OpenStudy (anonymous):

write the following sequence in an form and find the value of the 15th term 4,9,14,19. What is the sum of the first 15 terms of the above sequence?

OpenStudy (mathteacher1729):

Opening question: What is the pattern in this sequence? In other words, how do we get the next term, and the next, and the next...

OpenStudy (anonymous):

add 5 would the equation be an=-1 + 5n?

OpenStudy (anonymous):

it is an arithmatic progresion.series.... a = 4, d= 5, n=15 sum of 15 terms S15= n/2[ 2a +(n-1)d] = 15/2[ 8+14*5] = 15/2[78] = 15*39 = 585

OpenStudy (anonymous):

the formula for sum is s of n= n(a sub 1 + a sub n) divided by 2

OpenStudy (anonymous):

15 th term = a+(n-1)d = 4 + 14*4 =60

OpenStudy (anonymous):

i got 59..

OpenStudy (anonymous):

sorry, it is 4 + 14*5 = 74

OpenStudy (anonymous):

ok and what is the equation of the sequence in a sub n form

OpenStudy (anonymous):

do you have aim?

OpenStudy (anonymous):

already written......... a = first term d= common difference n= number of terms

OpenStudy (anonymous):

no it has to be like an=-1+5n is that right?

OpenStudy (anonymous):

for nth term a+(n-1)d for sum of first n terms n/2[2a +(n-1)d]

OpenStudy (anonymous):

yeah iknow but you plug in a as 4 and d as 5 and simplify what you can i thought?

OpenStudy (anonymous):

if you have problem to understand it , you r welcome to talk to me using skype, i m unable to explain it.......... skype id : sudhanshu_kmr

OpenStudy (anonymous):

do you have instant messenger?

OpenStudy (anonymous):

no......

OpenStudy (anonymous):

ok another question for you:\[\sum_{n=1}^{3}i^n-1\]

OpenStudy (anonymous):

its i to the n-1 power

OpenStudy (anonymous):

i +i^2 +i^3 -1 = i -1 - i -1= -2 as i^2 = -1, i^3 = -i

OpenStudy (anonymous):

i ended up with an answer of i?